A 2.71-g sample of benzene (C6H6) is burned in a constant-volume calorimeter that has a heat capacity of 11.8 kJ/°C. The temperature increases from 20.1 °C to 29.7 °C.

a) Determine the amount of heat released (in kJ).



kJ

b) Determine the molar energy of combustion of benzene (in kJ/mol).



kJ/mol
1
Expert's answer
2016-11-23T13:58:15-0500
1) The heat released in the calorimeter is Q = C*(t2 - t1) = 11.8 kJ/C * (29.7 C - 20.1 C) = 113 kJ.
2) The molar heat of combustion for C6H6 is H = Q*M(C6H6)/m(C6H6) = 113 kJ * 78 g/mol / 2.71 g = 3.25 kJ/mol.